A man had a sum of money which he sells among three children. The first took half of the money, the second took 2/3 of what he gave to the first. He the remainder to the third child. If the third removed Ghc5000, how much money did the man shared?
Answer by Theo(13342)

A man had a sum of money which he sells among three children. The first took half of the money, the second took 2/3 of what he gave to the first. He the remainder to the third child. If the third removed Ghc5000, how much money did the man shared?

let x = the sum of money.
he gave half the money to the first child.
this means he gave 1/2 * x to the first child and he had 1/2 * x left.
he gave 2/3 of what he gave to the first child to the second child.
this means he gave 2/3 * 1/3 * x to the second child = 2/6 * x to the second child.
this means he had 1/2 * x - 2/6 * x left which is equal to 3/6 * x - 2/6 * x which is equal to 1/6 * x.
he is left with 1/6 * x which he gave to the third child.
the third child received 5000.
1/6 * x = 5000.
x = 6 * 5000 = 30000.

let's see if the numbers check out.
he gave half of 30,000 to the first the first child which means he gave 15,000 to the first child and was left with 30,000 - 15,000 = 15,000.
he gave 2/3 of what he gave to the first child to the second child which means he gave 2/3 * 15000 to the second child = 10,000.
he was left with 15,000 - 10,000 = 5,000
he gave this to the third child.

the numbers check out and so the solution is assumed to be good.

the following table summarizes the transactions:


the man          his first child       his second child     his third child.

he starts with:

30,000

he give half of what he has to his first child:

15,000           15,000

he gives two thirds of what he has left to his second child:

5,000            15,000                10,000

he gives the rest to his third child:

0                15,000                10,000               5,000